OpenStudy (nathaliama13):

Which of the following ordered pairs represents the solution to the system given below? x − 3y = 12 2x + 6y = 60 (0, −4) (12, 0) (3, 21) (21, 3)

OpenStudy (nottim):

rearrange one of the equations to solve for y or x, and then sub it into another.

OpenStudy (anonymous):

So take the first equation and make it \[-3y = -x + 12\]

OpenStudy (anonymous):

Which is \[-y = \frac{ -x }{ -3} - 4\]

OpenStudy (anonymous):

i meant y not -y

OpenStudy (anonymous):

y = \[1/3x -4\]

OpenStudy (anonymous):

so then take that and plug it in to 2x + 6y = 60 \[2x + 6(1/3x−4) = 60\] and then simplify

OpenStudy (nottim):

Agh, don't tell them the answer.

OpenStudy (anonymous):

I didn't...

OpenStudy (nathaliama13):

@NotTim @Jackmullen10 do I use the substitution method.

OpenStudy (nottim):

Yes. Definitely. that's what he and I said.